Carrie Tune will receive $26,100 a year for the next 18 years as payment for a song she has just written. If a present 12 percent discount rate is applied: Use Appendix D. (Round "PV Factor" to 3 decimal places. Round the final answers to the nearest whole dollar.)
a-1. Calculate the present value.
Present value $ 189225
a-2. Should she be willing to sell out her future rights now for $178,000?
No
b-1. Calculate the present value if the payments will be made at the beginning of each year.
Present value $
b-2. Would she be willing to sell her future rights now for $178,000, if the payments will be made at the beginning of each year?
Answer #1

a-1)PMT =26100
Number of Years =18
Rate =12%
PV =PMT*((1-(1+r)^-n)/r) =26100*((1-(1+12%)^-18)/12%) =26100*7.250 =189225

a-2) NO you won't be selling

b-1) At beginning of year
PV =(1+r)*PMT*((1-(1+r)^-n)/r) =(1+12%)*26100*((1-(1+12%)^-18)/12%) =26100*8.120=211932

b-2) No , you won't be able to sell it